# MORPH: Shape-agnostic PDE Foundation Models

### v2

@@ 4 Experiments
4.1 Datasets
We select datasets from three publicly available PDE benchmarks i.e., PDEBENCH (Takamoto et al., 2022), pdegym Herde et al. (2024) and The Well (Ohana et al., 2024). They cover a wide variety of PDEs, the latter, however, focuses on more complex multiphysics problems. Because MORPH accommodates data heterogeneity, we demonstrate versatility in our dataset selection for both pretraining and fine-tuning. The datasets and their shapes are listed in Table 1. In the table, we display all fields in their inflated shapes separately if the dataset contains both scalar and vector fields. The six pretraining datasets are 1D Compressible Navier–Stokes (1D-CFD), 2D Incompressible Navier–Stokes (2D-CFD-IC), 2D Diffusion–Reaction (2D-DR), 2D Shallow-Water equations (2D-SW), 3D Compressible Navier–Stokes (3D-CFD) and 3D MagnetoHydroDynamics (3D-MHD). The seven fine-tuning datasets are 1D Diffusion–Reaction, 1D Burgers’ Equation (1D-BE), 2D Compressible Navier–Stokes (2D-CFD), 2D Gray–Scott Diffusion–Reaction, 2D Kolmogorov Flow (2D-FNS-KF), 3D Turbulent Gravity Cooling (3D-TGC), and 3D Compressible Navier–Stokes with Turbulence (3D-CFD-Turb). Compared to existing PDE foundation models, our pretraining and fine-tuning sets are larger and more diverse. We defer more dataset-specific details to Appendix A.
Table 1: Pretraining and finetuning (bold) datasets and their shapes. (N,T,F,C,D,H,W) denotes N = trajectories, T = time-steps, F = fields, C = components, D = depth, H = height, W = width. The datasets cover a wider range of heterogeneity in terms of various fields, their scalar or vector components and spatiotemporal resolutions. If the dataset contains both scalar and vector fields, we display all fields in their inflated shapes separately.
1D-CFD 1D-DR 1D-BE 2D-DR 2D-FNS-KF
(N,T,F,W) (N,T,W) (N,T,W) (N,T,F,H,W) (N,T,C,H,W)
(N,100,3,1024) (N,200,1024) (N,200,1024) (N,100,2,128,128) VV:(N,21,2,128,128)
2D-SW 2D-CFD 2D-CFD-IC 2D-GSDR
(N,T,H,W) (N,T,F,C,H,W) (N,T,C,H,W) (N,T,F,H,W)
(N,100,128,128) $V$: (N,21,1,2,512,512) $V$: (N,1000,2,512,512) (N,1001,2,128,128)
$\rho$: (N,21,1,1,512,512)
$P$: (N,21,1,1,512,512)
3D-MHD 3D-CFD 3D-CFD-Turb 3D-TGC
(N,T,F,C,D,H,W) (N,T,F,C,D,H,W) (N,T,F,C,D,H,W) (N,T,F,C,D,H,W)
VV:(N,100,1,3,64,64,64) VV:(N,21,1,3,128,128,128) VV:(N,21,1,3,64,64,64) VV:(N,T,1,3,64,64,64)
BB:(N,100,1,3,64,64,64) $\rho$:(N,21,1,1,128,128,128) $\rho$:(N,21,1,1,64,64,64) $\rho$:(N,T,1,1,64,64,64)
$\rho$:(N,100,1,1,64,64,64) PP:(N,21,1,1,128,128,128) PP:(N,21,1,1,64,64,64) PP:(N,T,1,1,64,64,64)
TT:(N,T,1,1,64,64,64)
4.2 Pretraining Setup
Balanced task sampling.
We train MORPH on multiple pretraining datasets with varying numbers of trajectories per dataset. We perform balanced task sampling to avoid imbalanced learning and catastrophic forgetting. Each dataset is assigned a sampling weight inversely proportional to its number of trajectories, so datasets with fewer trajectories are randomly sampled more frequently. In our corpus, the 3D-MHD dataset has the fewest trajectories and is therefore sampled more often than the remaining five. The resulting empirical per-batch sampling probabilities are approximately: 3D-MHD: 0.31, 2D-CFD-IC: 0.19, 3D-CFD: 0.18, 2D-DR: 0.12, 2D-SW: 0.12, and 1D-CFD: 0.08.
LoRA layers.
Low-Rank Adaptation (LoRA) is used for fine-tuning LLMs where the pretrained model weights are frozen while introducing trainable rank decomposition matrices into the transformer layers, dramatically reducing task-specific parameters during finetuning (Hu et al., 2022). The motivation is supported by evidence that over-parameterized networks possess low intrinsic dimensionality, suggesting that task-specific updates lie in a low-rank subspace (Aghajanyan et al., 2020). Mathematically, for a frozen weight $W_{0}\in\mathbb{R}^{d\times k}$, LoRA parameterizes the update as
yielding the adapted weight
$W^{\prime}=W_{0}+\frac{\alpha}{r}\,BA,$
Compute infrastructure.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@ 4.2 Pretraining Setup
Our model accommodates both fixed–order autoregressions, $AR(p)$, and time-varying orders, $AR(p_{1:T})$, where $p$ (or $p_{t}$) denotes the autoregressive order. In this paper, we train MORPH in a self-supervised $AR(1)$ setting where we learn a map $F_{\theta}:\mathcal{X}\to\mathcal{X}$ such that for all $t\geq 0$, $X_{t+1}\approx F_{\theta}(X_{t})$ by minimizing the mean-squared error. High–order fixed AR models (e.g., $AR(16)$ in McCabe et al. (2024) and $AR(10)$ in Hao et al. (2024)) often yield stronger long-horizon rollouts by reducing exposure bias via a longer receptive field. However, this design departs from the initial-value-problem (IVP) semantics that PDE foundation models aim to respect. For these reasons we adopt a one-step ($AR(1)$) formulation aligned with IVP semantics and preserves the compositional structure fundamental to PDE time evolution. Nevertheless, MORPH remains flexible and supports multi-order (fixed or varying) autoregression to improve rollout quality. Full pretraining and finetuning settings are documented in Appendix B.3,B.4.
@@ 4.3 Baselines
We evaluate MORPH against established baselines like Fourier Neural Operator (FNO) (Li et al., 2020) and U-Net (Ronneberger et al., 2015) as well as recent state-of-the-art foundation models: MPP (McCabe et al., 2024), DPOT (Hao et al., 2024), and Poseidon (Herde et al., 2024). FNO is a parameter-efficient neural operator that learns mappings between function spaces by approximating integral operators in Fourier space. It has become a popular choice for spatiotemporal surrogate modeling in fluid dynamics (Li et al., 2020). U-Net, originally proposed for biomedical image segmentation, is an encoder–decoder convolutional architecture with skip connections. Its strong multiscale feature extraction has led to widespread adoption in PDE surrogate modeling and operator learning (Rahman et al., 2022). MPP (McCabe et al., 2024) is among the first foundation models for PDEs, systematically demonstrating that multi-physics pretraining can outperform training from scratch on fine-tuned downstream tasks. DPOT (Hao et al., 2024) is a denoising pretraining operator transformer built around a Fourier-attention mixer and a temporal aggregation layer, trained with an autoregressive denoising objective. Poseidon (Herde et al., 2024) is an efficient multiscale operator transformer trained with an all-to-all scheme. It employs the scOT, a hierarchical multiscale transformer with shifted-window (Swin) attention and a lead-time–conditioned LayerNorm to support queries at continuous times.
@@
We follow the conventions of PDEBench (PB) and The Well (TW) when reporting errors on spatiotemporal PDE datasets. The Normalized Root Mean Squared Error (NRMSE) is the RMSE normalized by the $\ell_{2}$ norm of the ground truth, making it a scale-independent (dimensionless) measure and the default metric in PDEBench Takamoto et al. (2022). In contrast, the Variance-Scaled RMSE (VRMSE) normalizes the RMSE by the target’s per-snapshot standard deviation (i.e., the square root of its variance), thereby comparing errors to the intrinsic variability of the field. The Well adopts VRMSE as a primary metric. By construction, $\mathrm{VRMSE}(\bar{u},u)=1$, so values >1>1 indicate doing worse than predicting the per-snapshot spatial mean, while values <1<1 indicate improvement over that baseline Ohana et al. (2024).
A principal challenge in comparing MORPH to prior foundation models is data dimensionality. The existing models are trained on 2D benchmarks and are consequently constrained to 2D inputs, whereas MORPH is shape-agnostic. MORPH gives a single architecture that handles 11D, 22D, and 33D spatiotemporal fields without re-architecting. To ensure fairness, we report head-to-head results only where the data dimensionality aligns with a given baseline i.e., table entries are marked “-” when a model cannot be evaluated due to such shape constraints.
A further challenge for fair comparison is the input context window employed by prior foundation models. MPP is trained with a $16$-step context, while DPOT uses a $10$-step context. In line with numerical solvers, we posit that PDE foundation models should initialize from an initial condition (or initial time step) and advance the state autoregressively into the future. Accordingly, we train with a single-step context, $\mathrm{AR}(1)$, for all models. Naturally, models trained with longer autoregressive contexts than $\mathrm{AR}(1)$ may enjoy an advantage on standard error metrics and better performance on rollouts. We therefore interpret cross-model comparisons with this caveat in mind and present the main results in the next subsection.
@@ 4.4 Main results
Table 2: NRMSE (N) and VRMSE (V) of the experiments (lower is better). The table is organized into two parts, each split into pretraining (PT) and fine-tuning (FT) sections: (a) task-specific standalone surrogate (SS) models trained on their respective datasets, (b) foundation models (FM) pretrained on PT sets (zero-shot results) and fine-tuned on FT sets (full-shot results). Within each section, the best score for the section is typeset in bold. When a score is the global best for a dataset, it is shown as bold+underline. Our standalone models are shown in blue text and foundation models in red. An asterisk () on the L model indicates LoRA fine-tuning.
@@
1D-DR (N) - - - - 0.0008 0.00131 0.00516 0.00225
1D-BE (N) - - - - 0.0302 0.0584 0.06605 0.06907
2D-CFD (N) - - 0.55897 0.59939 0.05401 0.05104 0.05886 0.0423
2D-GSDR (V) 0.00749 0.00744 0.01199 0.01416 0.00763 0.00725 0.00757 0.01098
3D-CFD-Turb (N) - - - - 0.10742 0.10205 0.07131 0.09942
3D-TGC (V) - - - - 0.05312 0.04136 0.04811 0.03946
Our primary experiments are organized in two parts or four sections in Table 2. We compare baseline and SOTA models using NRMSE and VRMSE as evaluation metrics. The top section reports standalone surrogate (SS) models (Ti, $\approx 7$M parameters; S, $\approx 30$M parameters) evaluated on both the pretraining and fine-tuning datasets. The bottom section reports foundation model (FM) results under zero-shot and full-shot fine-tuning settings. In the top section, we evaluate MORPH on standalone pretraining and finetuning sets. Our MORPH-SS-Ti and S models outperform baseline and SOTA models on standalone datasets and remain competitive across the rest. Practically, this makes MORPH a viable standalone surrogate when the target application focuses on a single PDE family. In the bottom section, we evaluate zero-shot performance on the pretrained sets and full-shot on finetuning sets against SOTA models. Zero-shot results assess cross-physics generalization and retention after multi-dataset pretraining. Deviations on pretrained datasets from standalone baselines indicate interference (e.g., gradual forgetting of task-specific biases). On the other hand, full-shot results quantify adaptation-efficiency gains over standalone surrogates on finetuning datasets.
MORPH Standalone vs foundation models.
We compare our standalone and foundation models under both zero-shot (pretraining-only evaluation) and full-shot (fine-tuning) regimes. MORPH-FM outperforms MORPH-SS on 3 out of 4 zero-shot tests and on 5 out of 6 fine-tuned benchmarks. Across these settings, our models also surpass strong baselines and existing models. The autoregressive rollout results for both standalone and fine-tuned models are deferred to Appendix C.2. On the FNS-KF prediction task (See Figs. 6 and 7), MORPH-FM-S achieves better rollouts than MORPH-FM-Ti, which is reflected in lower MSE as well (0.00089 for S vs. 0.00473 for Ti).
@@
We investigate fine-tuning the MORPH-FM-L model using low-rank adapters (LoRA). In our setup, only 77M of the 480M parameters comprising LoRA adapters on attention and MLP blocks, positional encodings and layer norms are fine-tuned, which is fewer trainable parameters than in the M model (126M). Notably, MORPH-FM-L matches or outperforms the M model across zero-shot evaluations and full-shot fine-tuning. To our knowledge, this is the first study to apply LoRA to PDE foundation models, and our results indicate that LoRA-based fine-tuning is a promising direction given the expanding PDE datasets and increasing model sizes. Future work will examine LoRA-based fine-tuning more extensively, including the effect of adapter rank across different datasets.
Data and compute-scarce finetuning.

### v3

@@ 4 Experiments
Datasets.
We select datasets from three publicly available PDE benchmarks, i.e., PDEBench (takamoto2022pdebench), PDEgym (herde2024poseidon) and The Well (ohana2024well). They cover a wide variety of PDEs; the latter, however, focuses on more complex multiphysics problems. Because MORPH accommodates data heterogeneity, we demonstrate versatility in our dataset selection for both pretraining and fine-tuning. The pretraining and finetuning datasets and their shapes are listed in Table 1 and 2. Compared to existing PDE foundation models, our pretraining and fine-tuning sets are larger and more diverse. We defer more dataset-specific details to Appendix A.
Table 1: Pretraining datasets with their physical fields (scalar or vector) and data shapes (N,T,D,H,W,C), together with the benchmark/source.
PT sets
MHD-3D
DR-2D
CFD-1D
CFD2D-IC
CFD-3D
SW-2D
Shapes
$\rho$: $(97,100,64^{3})$
$\mathbf{B}$: $(97,100,64^{3},3)$
$\mathbf{v}$: $(97,100,64^{3},3)$
$u_{1}$: $(1k,100,128^{2})$
$u_{2}$: $(1k,100,128^{2})$
$\rho$: $(10k,100,1024)$
$\mathbf{v}$: $(10k,100,1024)$
$P$: $(10k,100,1024)$
$\rho$: $(64,1k,512^{2})$
$\mathbf{v}$: $(64,1k,512^{2},2)$
$P$: $(64,1k,512^{2})$
$\rho$: $(200,21,128^{3})$
$\mathbf{v}$: $(200,21,128^{3},3)$
$P$: $(200,21,128^{3})$
$h$: $(1k,100,128^{2})$
Source The Well PDEBench PDEBench PDEBench PDEBench PDEBench
Table 2: Finetuning datasets with their physical fields (scalar or vector) and data shapes (N,T,D,H,W,C), together with the benchmark/source.
FT sets
DR-1D
CFD-2D
CFD3D-TURB
BE-1D
GSDR-2D
TGC-3D
FNS-KF-2D
Shapes
$u_{1}$: $(10k,100,1024)$
$\rho$: $(10k,21,512^{2})$
$\mathbf{v}$: $(10k,21,512^{2},2)$
$P$: $(10k,21,512^{2})$
$\rho$: $(600,21,64^{3})$
$\mathbf{v}$: $(600,21,64^{3},3)$
$P$: $(600,21,64^{3})$
$u$: $(10k,200,1024)$
AA: $(200,1k,128^{2})$
$B$: $(200,1k,128^{2})$
$\rho$: $(100,50,64^{3})$
$\mathbf{v}$: $(100,50,64^{3},3)$
$P$: $(100,50,64^{3})$
$T$: $(100,50,64^{3})$
$u$: $(20k,21,128^{2})$
$v$: $(20k,21,128^{2})$
Source PDEBench PDEBench PDEBench PDEBench The Well The Well PDEgym
@@
Our model accommodates both fixed–order autoregressions, $AR(p)$, and time-varying orders, $AR(p_{1:T})$, where $p$ (or $p_{t}$) denotes the autoregressive order. In this paper, we train MORPH in a self-supervised $AR(1)$ setting where we learn a map $F_{\theta}:\mathcal{X}\to\mathcal{X}$ such that for all $t\geq 0$, $X_{t+1}\approx F_{\theta}(X_{t})$ by minimizing the mean-squared error. High–order fixed AR models (e.g., $AR(16)$ in mccabe2024multiple and $AR(10)$ in hao2024dpot) often yield stronger long-horizon rollouts by reducing exposure bias via a longer receptive field. However, this design departs from the initial-value-problem (IVP) semantics that PDE foundation models aim to respect. For these reasons we adopt a one-step ($AR(1)$) formulation aligned with IVP semantics that preserves the compositional structure fundamental to PDE time evolution. Nevertheless, MORPH remains flexible and supports multi-order (fixed or varying) autoregression to improve rollout quality. For brevity, full pretraining and finetuning settings are documented in Appendix B.2, B.3.
@@ 4.1 Baselines
We compare MORPH against classical neural operators and recent PDE foundation models. As popular baselines, we use Fourier Neural Operator (FNO) (li2020fourier), a parameter-efficient spectral neural operator widely adopted for fluid surrogates, and UNet (ronneberger2015u; rahman2022u), a convolutional encoder–decoder with multiscale skip connections. Among foundation models, we include MPP (mccabe2024multiple), which demonstrates gains from multi-physics pretraining; DPOT (hao2024dpot), a denoising operator transformer with Fourier attention; and Poseidon (herde2024poseidon), a hierarchical multiscale operator transformer with shifted-window attention and lead-time-conditioned normalization for continuous-time querying.
@@
We follow the conventions of PDEBench (PB) and The Well (TW) when reporting errors on spatiotemporal PDE datasets. The Normalized Root Mean Squared Error (NRMSE) is the RMSE normalized by the $\ell_{2}$ norm of the ground truth, making it a scale-independent (dimensionless) measure and the default metric in PDEBench (takamoto2022pdebench). In contrast, the Variance-Scaled RMSE (VRMSE) normalizes the RMSE by the target’s per-snapshot standard deviation (i.e., the square root of its variance), thereby comparing errors to the intrinsic variability of the field. The Well adopts VRMSE as a primary metric (ohana2024well).
Comparison Caveats.
A central challenge in fairly comparing PDE foundation models is their use of different pretraining corpora. In this work, we therefore fine-tune the FMs on the same downstream tasks and interpret the results independently of their pretraining corpus. Another difficulty in comparing MORPH to prior foundation models is the data modality. The existing models are trained on 2D benchmarks and are consequently constrained to 2D inputs, whereas MORPH is agnostic to data modality. To ensure fairness, we report head-to-head results only where the data modality aligns with a given baseline, i.e., table entries are marked “-” when a model cannot be evaluated due to such shape constraints. A further complication for fair comparison is the input context window employed by prior foundation models. MPP is trained with a $16$-step context, while DPOT uses a $10$-step context. In line with numerical solvers, we posit that PDE foundation models should initialize from an initial condition (or initial time step) and advance the state autoregressively into the future. Accordingly, we train with a single-step context, $\mathrm{AR}(1)$, for all models.
@@ 4.2 Main results
Our primary experiments are organized in Tables 3,4,5. In these tables, we compare baseline and SOTA models using NRMSE and VRMSE as evaluation metrics.
Table 3: Standalone surrogates: NRMSE (N) and VRMSE (V) of the experiments (lower is better) on pretraining and finetuning datasets. The best score is typeset in bold. The global best for the dataset is bold + underlined. Our standalone models are shown in blue text.
@@
Table 4: Zero-shot performance of foundation models: NRMSE (N) and VRMSE (V) of the experiments (lower is better). The best score is typeset in bold. The global best for the dataset is bold + underlined. Our foundation models in red.
Zero-shot on pretraining sets Zero-shot on finetuning sets
Datasets Models Ti (7M) S (30M) M (126M) L (480M) Datasets Models Ti (7M) S (30M) M (126M) L (480M)
1D-CFD (N) 0.04203 0.05715 0.06534 0.05056 1D-DR (N) 0.02147 0.02004 0.01951 0.02553
2D-DR (N) 0.12588 0.11843 0.13524 0.11067 1D-BE (N) 0.20365 0.21151 0.20735 0.20164
2D-CFD-IC (N) 0.09657 0.13037 0.09509 0.08567 2D-CFD (N) 0.10084 0.10066 0.09988 0.09019
2D-SW (N) 0.00751 0.00605 0.00603 0.00454 2D-GSDR (V) 0.50432 0.51025 0.49516 0.48519
3D-MHD (V) 0.32203 0.31276 0.32156 0.28496 3D-CFD-Turb (N) 0.31692 0.29181 0.29507 0.28436
3D-CFD (N) 0.09163 0.11308 0.11541 0.07279 3D-TGC (V) 0.53468 0.48879 0.49386 0.47722
Table 5: Full-shot finetuning results for foundation models: NRMSE (N) and VRMSE (V) of the experiments (lower is better). The best score is typeset in bold. The global best for the dataset is bold + underlined. Our foundation models in red. An asterisk () on the L model indicates LoRA fine-tuning.
@@
1D-DR (N) - - - - 0.0008 0.00131 0.00173 0.00135
1D-BE (N) - - - - 0.0302 0.0584 0.04177 0.04123
2D-CFD (N) 0.77357 0.63255 0.55897 0.59939 0.05401 0.05104 0.05886 0.0423
2D-GSDR (V) 0.00749 0.00744 0.01199 0.01416 0.00763 0.00725 0.00518 0.00497
3D-CFD-Turb (N) - - - - 0.10742 0.10205 0.08635 0.09170
3D-TGC (V) - - - - 0.05312 0.04136 0.03756 0.03246
Standalone surrogates.
In Table 3, we compare performance of MORPH standalone surrogates (MORPH-SS) on pretraining and finetuning datasets against existing baseline (FNO and UNets) and SOTA models (MPP and DPOT). The results for baselines and SOTA models are inherited from the respective benchmark sources (takamoto2022pdebench; ohana2024well) and articles (mccabe2024multiple; hao2024dpot). Our MORPH-SS-Ti and S models outperform baseline and SOTA models on standalone datasets and remain competitive across the rest. Practically, this makes MORPH a viable standalone surrogate when the target application focuses on a single PDE family.
Zero-shot performance.
In Table 4, we report the zero-shot performance of different MORPH foundation models. For this evaluation, we run the pretrained FMs in inference mode on the test sets of their corresponding datasets. We observe that MORPH-FM-L achieves the best performance on 11 out of 12 datasets. These results indicate favorable scaling behavior of MORPH under zero-shot adaptation.
Full-shot finetuning performance.
Full-shot fine-tuning results are presented in Table 5. On the 2D datasets, we compare the fine-tuning performance of MORPH-FM to DPOT-FM and Poseidon-FM, and find that our models consistently surpass these existing FMs. Together with Tables 3 and 5, we can quantify the adaptation-efficiency gains of MORPH-FM over MORPH-SS on the fine-tuning benchmarks: MORPH-FM outperforms MORPH-SS on 5 out of 6 fine-tuning datasets.
@@
We investigate fine-tuning the MORPH-FM-L model using low-rank adapters (LoRA). In our setup, only 77M of the 480M parameters comprising LoRA adapters on attention and MLP blocks, positional encodings and layer norms are fine-tuned, which is fewer trainable parameters than in the M model (126M). Notably, MORPH-FM-L matches or outperforms the M model across full-shot fine-tuning (Table 5). To our knowledge, this is one of the first two concurrent studies to apply LoRA to PDE foundation models (see also: holzschuh2025pde), and our results indicate that LoRA-based fine-tuning is a promising direction given the expanding PDE datasets and increasing model sizes. Future work will examine LoRA-based fine-tuning more extensively, including the effect of adapter rank across different datasets.
Autoregressive rollouts.
We have presented autoregressive rollouts results for standalone model and foundation model. Figs. 6 and 7 (Appendix C.3) show rollout results for MORPH-SS-Ti and MORPH-SS-S models on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 3. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon. Fig. 3 presents autoregressive rollouts for MORPH-FM-S fine-tuned on the FNS-KF prediction task for $v_{y}$ ($v_{x}$ in Appendix C.3 Fig 9). It can be observed that MORPH-FM-S achieves better rollouts than MORPH-FM-Ti (Fig. 8 in Appendix C.3).
Figure 3: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts of $v_{y}$ with $t=0$ (initial frame) as input.
4.3 Scaling studies
We study scaling behavior in both pretraining and fine-tuning, with respect to both dataset size and model capacity. For pretraining, Fig. 5 (Appendix C.1) reports training and validation losses for MORPH-FM-S (30M parameters) as we vary the number of pretraining trajectories. The results are shown over the first 10 epochs, during which both losses decrease steadily as the pretraining dataset size increases. For brevity, more details are deferred to Appendix C.1. Turning to model capacity, our largest model (MORPH-FM-L) ultimately saturates at a validation loss of $\sim$0.02 during pretraining, compared to $\sim$0.09 for the Ti/S variants. However, for a fixed training budget of 10 epochs under a single shared protocol (same learning rate, batch size, and weight decay), we do not observe clear gains when scaling from Ti to L. This suggests that mild model-specific hyperparameter tuning and/or extended training is needed for the increased capacity to operate in the regime where performance scales predictably with model size.
Scaling w.r.t. finetuning dataset size.
Fig. 4 presents a dataset–size scaling study on two downstream tasks (1D-DR and 2D-FNS-KF). We fix the training protocol (optimizer, schedule, epochs, batch size) and vary the number of fine-tuning trajectories. We report RMSE as a function of the number of trajectories used to fine-tune MORPH-FM, and compare against the standalone MORPH-SS trained on 100% of trajectories. Errors consistently decrease as dataset size increases. On 1D-DR, the pretrained MORPH-FM-Ti already outperforms the standalone model with only 25% of the trajectories. On 2D-FNS-KF, with $\sim$ 256 trajectories (less than 1% of the full set), MORPH-FM-S surpasses the standalone MORPH-SS-S trained on all trajectories.
Figure 4: Scaling with respect to finetuning dataset size: (Left:) MORPH-FM-Ti on 1D-DR: RMSE vs $\%$ different trajectories finetuned for 100 epochs, (Right:) MORPH-FM-S on 2D-FNS-KF: RMSE vs # trajectories of finetuend for 100 epochs.
Scaling w.r.t. finetuning model size (across FMs).